OpenStudy (anonymous):

what is the equation of the line that passes through the points (0,4) and (3,-2) A.y=2x+0 B.y=- 1/2x+ (-2) C.y=-2x+4 D.y=1/2x+4 my answer is b is that correct ?

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

help

OpenStudy (trojanpoem):

dy/dx = y2- y1/x2-x1 = slope -2 - 4 / 3 - 0 -6/3 = -2 y = mx + b ( m = slope) y = -2 x + b (put (0,4) ) 4 = b , b = 4 y = -2x + 4

OpenStudy (anonymous):

oh

OpenStudy (trojanpoem):

Show me your steps.

OpenStudy (anonymous):

k